Monochromatic light passes through two slits separated by a distance of 0.0340 mm .If the angle to the third maximum above the central fringe is 3.31 ∘ , what is the wavelength of the light? =_______nm
Answer #1

Given that

Monochromatic light passes through two slits separated by adistanceof ( 2d) = 0.0340 mm.

=0.034*10-3m

The angle to the third maximum above the central fringe is(θ) =3.31°

Then the wavelength of the light is gvien by

sinθ=nλ/2d

Then the wavelength of the light is

λ=2dsinθ/n where n =3

Now substitute all the above values you get therequiredanswer

λ=(0.034*10-3)sin3.31/3

λ=654.37nm